From nch@roe.ac.uk Fri Nov 30 15:23:38 2012 Date: Fri, 30 Nov 2012 14:13:18 +0000 From: Nigel Hambly To: Eckhard Sutorius , Nicholas Cross , Rob Blake , Ross Collins , Mike Read , Clive Davenhall , Mark Holliman Cc: Lorenzo Rimoldini , Bob Mann , Jim Emerson , Keith Noddle , Norman Gray , Tom Shanks , Stephen Warren , Simon Dye , Stelios Voutsinas , Tom Kerr , Mike Irwin , Andy Lawrence at ROE , Nigel Metcalfe Subject: WFAU VDFS Science Archives project team meeting minutes, 30 Nov 2012 Present: MAR, RSC, RPB, ETWS, NJC, NCH, RGM Apologies: MSH, JPE, AL, STV, NG, KTN DoNM: 10am Friday December 14th 2012 *** NB in the Villas meeting room *** MAR in the chair (room is booked). Actions discharged this time: ----------------------------- ACTION: ETWS/MAR to check disk38 catalogues for corruption Discharged; no major problems found ACTION: NJC to find out where we stand with regard to a VIKING ESO release. Discharged; a P86 release is required, consisting of ~150 pointings. The processing is starting today (see below). ACTION: MAR to check which semesters are to be included in the next VVV release. Discharged; P87 is now required. Actions partly discharged but continuing: ----------------------------------------- ACTION: ETWS (with help from NJC as necessary) to add FeII 1.644 micron filter details into the WSA. Continues; all done apart from Vega-to-AB and zeropoint attributes (the later needs the latest data to be ingested). Actions carried forward from 16/11/12 meeting --------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. Specific points and new actions: -------------------------------- Project management: Nothing for the minutes. WFCAM, VISTA and VST updates: It was noted that feedback from ESO following survey progress reviews has been generally positive for those that we've been CC'd on (VMC and VVV). Comments and issues arising from CASU minutes: No new minutes this time. Networking: ETWS noted that transfers are now up-to-date as at the end of August 2012. Transfer speeds holding up at healthy 30 (OSA), 20 (VSA) and 5 (WSA) MB/s. ETWS also noted that VMC have asked for the Japanese AKARI mid-IR catalogues to be imported into the VSA. It was pointed out that this is OK provided the size of the catalogue warrants inclusion, otherwise crossmatches etc. should be done "client-side" (e.g. using topcat). WSA/VSA/OSA/GES Operations: ETWS reported: -- Transferred the OSA data up to and including September 2012, the VSA up to August 2012, and the WSA up to August 2012. -- The OSA data is also nearly completely ingested apart from catalogues for September. -- Updated the code to speed up VVV ingests by moving the update of objIDs to the FinaliseCatalogueIngest script, which is run after all ingests have taken place. This reduced the ingest time for P88 from 29 days to 4 days. -- Finished fixing all the disk38 pixel data for the VSA, the catalogue data shows no problems (yet). Nearly finished recreating all the JPGs that were on disk38, since there was no easy way to check for broken ones. RPB noted that VVV tile-pawprint operations have been speeded up significantly by attaching further DB constraints to the monthly partitioned tables. Hardware and Systems: The disk38 saga is drawing to a close - the team decided to give up on the slight possibility of recovering data from the constituent disks and have returned them to the vendors with a firmly worded "replace these please." 12x4TB devices have been offered in exchange for the 24x2TB originals; our friendly suppliers are going to have a go at testing the disks but we suspect they'll be unsuccessful given all the tests we've done. Survey Data Release: UKIDSS DR10 - QC1 is being closed out today. Next up is quality bit flagging, and (for the DXS) checking that images on the infamous disk38 are OK for deep stacking. ACTION: RPB to set in motion q-bits for LAS/GCS once given the nod by MAR ACTION: ETWS to check the DXS fileset (provided by RPB) for corruption NJC is looking after the necessary steps to incorporate the recently delivered UDS; it was noted that the GPS will not be included in the initial DR10 release as eyeballing and the usual small amount of reprocessing is yet to be completed. It was noted that WISE-ASS (that's All Sky Survey for the uninitiated) crossmatches should be set up for UKIDSS, since presently the crossmatch definitions point to the old preliminary source catalogue. ACTION: RPB to set up WISE-ASS cross-neighbours for UKIDSS surveys to replace the old preliminary source catalogue matchings VISTA: a new consortium release (incorporating observations up to P88) has been produced, along with the associated ESO deliverables that also include P88 deep stacks). A consortium release DB of VIKING is currently being created. VVV consortium release is currently running at the tile-pawprint stage. ATLAS - some response from Durham concerning QC questions; MAR will continue to liaise with them over a first release. Software: RSC noted: "The detection quality flagging code has been updated to support the new monthly-split detection tables. During testing of this it was noticed that a handful of VMC, VHS, VIDEO and probably VVV tiles were missing the bit 12 flag (Source in tile region of detector 16). This will be fixed up for future releases. Still looking into an improved saturation flag for the VVV detections." MAR noted some MultiGetImage updates on the VSA UI as requested by one of our most loyal customers. NJC has been making a few small mods and fixes during investigating the photometry issues reported last time. Non-survey Data Release: NCH asked about VSA PI programmes, and ETWS noted that P89 ingests are next in the operations queue after which it will be possible to set up some DBs (PIs for the use of). Astrogrid deployment, VO & Data Analysis services: It was brought to our attention (via a somewhat curious route) that many of WFAU's VO cone-search services were offline earlier in the week. MSH has fixed them, noting that they seem to have been brought to their knees by somebody hammering them incessantly and possibly with a faulty script. Anyway, the situation is being monitored ... Miscellaneous: ... in between nappy changes, feedings and generally sleepless nights on the part of MSH (WFAU extends hearty congrats to you and Gila on the birth of baby Maya). RGM noted that STV and DM will be around next week; NJC noted that the VSA paper is going to appear (in the bumper Xmas issue?) in A&A. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. ============================================================= -- Scanned by iCritical.